# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ 6.2 The order of the asymmetry operator
Definition 6.1 implies that, a priori, our asymmetry operator $A$ is a pseudodifferential operator of order 0. The goal of this subsection is to show that, as a result of cancellations in the symbol, the order of $A$ is in fact $-3$. This will be done in three steps.
@@
    First, we will show that $A$ is of order $-2$. This will follow easily from the facts already established earlier in the paper.
@@
    Next, we will show that, in normal coordinates, parallel transport does not contribute to homogeneous components of the symbol of $A$ of degree $-2$ and $-3$.
@@
    Finally, we will show that $A$ is of order $-3$. This is the subject of Theorem 6.6 which will come in the end of the subsection and is one of the main results of this paper.
@@
The asymmetry operator $A$ is a pseudodifferential operator of order $-2$.
@@
$[p_{\pm}]_{\alpha}{}^{\beta}(x,\xi)\sim\sum_{j=0}^{+\infty}[(p_{\pm})_{-j}]_{\alpha}{}^{\beta}(x,\xi),\qquad[(p_{\pm})_{-k}]_{\alpha}{}^{\beta}(x,\lambda\xi)=\lambda^{-k}[(p_{\pm})_{-k}]_{\alpha}{}^{\beta}(x,\xi)\quad\forall\lambda>0,$ (6.5)
@@
The decomposition (6.6) is not invariant, it relies on our particular choice of local coordinates but it is convenient for our purposes. Namely, in what follows we will show that $A_{\mathrm{pt}}$ will not contribute to the principal symbol of order $-3$ of the asymmetry operator $A$ at the point $z$.
@@
In what follows, we will be using the expansion (E.8) for $Z_{\alpha}{}^{\beta}$.
The claim (6.9) for $j=0$ and $j=1$ follows immediately from (E.8).
@@
Upon integration by parts, we get from (E.8)
@@
where ||$\cdot$|| ⋅ | is the Euclidean norm. By the symmetries of the Riemann tensor we have
$\operatorname{Riem}_{\alpha\mu\kappa\nu}=\frac{1}{2}(R_{\alpha\mu\kappa\nu}+\operatorname{Riem}_{\kappa\nu\alpha\mu}).$ (6.11)
@@
The expansion (E.8) and formula (6.14) imply, upon integration by parts,
@@
The asymmetry operator $A$ is a pseudodifferential operator of order $-3$.
@@
The principal symbol $A_{\mathrm{prin},2}$ (that is, the principal symbol of the asymmetry operator $A$ viewed as an operator of order $-2$) should be proportional to
@@
$[(a_{\mathrm{diag}})_{-2}](z,\xi)=[(p_{+}-p_{-})_{-2}]_{\alpha}{}^{\alpha}(z,\xi)=0.$ (6.22)
@@
Simplification 3. It is clear that $\operatorname{Ric}$ will appear in $[(a_{\mathrm{diag}})_{-2}](z,\xi)=[(p_{+}-p_{-})_{-2}]_{\alpha}{}^{\alpha}(0,\xi)$ in a linear fashion. This means that we can set
@@
  • one can scale $\xi$ so as to normalise it, using the homogeneity of $[(a_{\mathrm{diag}})_{-2}](0,\xi)$.
@@
Simplification 5. In implementing the algorithm, expand all quantities at the $k$-th step in powers of $x$ and $\eta$ retaining only terms up to order $2-k$, jointly in $x$ and $\eta$. Here we are assuming that $x$ and $\eta$ are of the same order.
Simplifications 1–5 reduce our algorithm to the manipulation of polynomials. Note that the vector spaces of polynomials in the variables $x^{\alpha}$ and $\eta_{\alpha}\,$, $\alpha=1,2,3$, of degree 0, 1 and 2 have dimension 1, 7 and 28 respectively [^4] The dimension of the vector space of polynomials of degree $k$ in $n$ variables is (${{k+n}\choose{k}}\,.$.
@@
$[\operatorname{tr}(X_{+,2}-X_{-,2})_{\mathrm{prin}}](0,\xi_{0})=0.$ (6.29)
@@
where $(P_{\pm,1}^{2})_{-2}$ stands for the homogeneous component of degree $-2$ of the symbol of the operator $P_{\pm,1}^{2}\,$. To compute the latter, one needs to use the formula for the symbol of the composition of pseudodifferential operators: given two operators $A$ and $B$ with symbols $a$ and $b$, the symbol $\sigma_{BA}$ of their composition $BA$ is given by the formula
@@ 6.3 The principal symbol of the asymmetry operator
The principal symbol $A_{\mathrm{prin},3}$ (that is, the principal symbol of the asymmetry operator $A$ viewed as an operator of order $-3$) should be proportional to
@@
$[(p_{+}-p_{-})_{-3}]_{\alpha}{}^{\alpha}(z,\xi)=-\frac{1}{2|\xi|^{5}}\varepsilon^{\alpha\beta\gamma}\,\nabla_{\alpha}\operatorname{Ric}_{\beta}{}^{\rho}(z)\,\xi_{\gamma}\xi_{\rho}\,.$ (6.32)
The strategy for proving (6.32) consists in implementing the algorithm from Proposition 5.3 in chosen normal coordinates up to order $-3$. Even with the help of modern computer algebra, doing this to the required accuracy is a highly nontrivial task. Therefore, as in the proof of Theorem 6.6, we describe below some helpful simplifications, which once again reduce all calculations to the manipulation of polynomials.
@@
Simplification 1. It is known, see [48, p. 211 formula (3.4)] or [45, formula (6)], that
@@
Simplification 5. In implementing the algorithm, expand all quantities at the $k$-th step in powers of $x$ and $\eta$ retaining only terms up order $3-k$, jointly in $x$ and $\eta$. Here we are assuming that $x$ and $\eta$ are of the same order.
@@
$[\operatorname{tr}(X_{+,3}-X_{-,3})_{\mathrm{prin}}](0,\xi_{0})=-\frac{1}{2}.$ (6.36)
@@
Since formula (1.13) is one of the key results of this paper, we provide in Appendix F an independent verification relying on formula (5.10).
@@ 7.1 Structure of the integral kernel near the diagonal
Now, the operator $A$ is not guaranteed to be of trace class, so one cannot apply formula (4.14) to it. The issue here is that in dimension 3 a sufficient condition for a pseudodifferential operator to be of trace class is that its order be strictly less than $-3$, whereas according to Theorem 1.4 for a generic Riemannian 3-manifold the order of $A$ is exactly $-3$. This means that we need to perform some sort of very basic regularisation in order to get over the finish line and define, in an invariant and geometrically meaningful manner, the trace of $A$
@@
This is an important property which will be used on two occasions in the subsequent arguments in this section. We will see that property (7.4) implies that the singularity of the integral kernel of our asymmetry operator $A$ is weaker than one would expect of a generic pseudodifferential operator of order $-3$. For a generic pseudodifferential operator of order $-3$ one would have a logarithm in the leading term of the singularity of the integral kernel, but there is no logarithm in (7.3).
@@
$\begin{split}\frac{1}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}\frac{1}{\langle\xi\rangle^{5}}\,e^{-iy^{\mu}\,\xi_{\mu}}\,\,\mathrm{d}\xi&=\frac{1}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}\frac{1}{\langle\xi\rangle^{5}}\,\cos\left(|y|\,\xi_{1}\right)\,\,\mathrm{d}\xi\,\\ &=\frac{2\pi}{(2\pi)^{3}}\int_{-\infty}^{+\infty}\cos\left(|y|\,\xi_{1}\right)\left(\int_{0}^{+\infty}\frac{r}{(1+(\xi_{1})^{2}+r^{2})^{5/2}}\,\mathrm{d}r\right)\,\mathrm{d}\xi_{1}\,\\ &=\frac{1}{12\pi^{2}}\int_{-\infty}^{+\infty}\frac{\cos\left(|y|\,\xi_{1}\right)}{(1+(\xi_{1})^{2})^{3/2}}\,\mathrm{d}\xi_{1}\,,\end{split}$ (7.7)
@@
[…]
The latter is known as Basset’s Integral [21, (10.32.11)]. ∎
@@
where $\gamma$ is the Euler–Mascheroni constant — see, e.g., [21, (10.31.11) and (10.25.2)]. By combining (7.9) and (7.6) one obtains (7.8). ∎
@@
In the previous section we have established that the asymmetry operator $A$ is a pseudodifferential operator of order $-3$ with principal symbol (1.13). Locally, the symbol $a(x,\xi)$ of $A$ and its integral kernel $\mathfrak{a}(x,y)$ are related via the (distributional) identity
@@
Since a pseudodifferential operator of order $-4$ on a 3-manifold has a continuous integral kernel, formulae (1.13), (1.10) and (7.11) imply that problem at hand reduces, in normal coordinates centred at $x$, to examining the quantity
$-\frac{1}{16\pi^{3}}\varepsilon^{\alpha\beta\gamma}\,\nabla_{\alpha}\operatorname{Ric}_{\beta}{}^{\rho}(0)\,\int_{\mathbb{R}^{3}}e^{-iy^{\mu}\,\xi_{\mu}}\,\frac{\xi_{\gamma}\xi_{\rho}}{|\xi|^{5}}\,(1-\chi(|\xi|))\,\mathrm{d}\xi\,,$
@@
so the required result follows from Lemma 7.5. Note that the term with $\delta_{\gamma\rho}$ from (7.10) vanishes upon contraction with $\,\varepsilon^{\alpha\beta\gamma}\,\nabla_{\alpha}\operatorname{Ric}_{\beta}{}^{\rho}(0)\,$ due to formula (7.4). ∎
@@ 7.2 Local and global trace
The issue at hand is that the expression (7.13) is not a scalar because the quantities $(x-y)^{\gamma}$ and $(x-y)^{\rho}$ are not vectors based at the point $x$. Let us switch from local coordinates $x$ to local coordinates $\widetilde{x}$ and let $\widetilde{f}(\widetilde{x},\widetilde{y})$ be the representation of the quantity (7.13) in these new coordinates. It is easy to see, by writing down the appropriate Jacobians, that
@@
Now, let $\widetilde{\mathfrak{a}}_{c}(\widetilde{x},\widetilde{y})$ be the representation of the quantity $\mathfrak{a}_{c}(x,y)$ in the new coordinates. Formulae (7.2), (7.3), (7.13), (7.14) and the fact that $\mathfrak{a}(x,y)$ is a true scalar in two variables imply
@@
[…]
@@ 8 Challenges in higher dimensions
One could have also considered the case $d=4k+1$, $k=1,2,\ldots$, with $\operatorname{curl}$ acting in $\Omega^{2k}(M)$ in accordance with (1.5). However, in this case $\operatorname{curl}$ is formally anti-self-adjoint, the spectrum is purely imaginary and there is no spectral asymmetry [7, Theorem 3.2]. As our interest is in spectral asymmetry, we do not discuss the case $d=4k+1$. Note also that the spectral problem for $\operatorname{curl}$ in dimension $d=4k+1$ cannot be formulated in terms of the spectral theory of self-adjoint operators in real Hilbert spaces: one can introduce the factor $\,i\,$ in the right-hand side of (1.5) to make the operator formally self-adjoint (as was done in [7, Definition 2.2]), but the resulting operator will no longer be an operator acting in the vector space of real differential forms. The bottom line is that the two cases, $d=4k+3$ and $d=4k+1$, are fundamentally different. We stick with (8.1) because this gives the only genuine generalisation of the operator $\operatorname{curl}$ to higher dimensions.
@@
First challenge. We expect that in higher dimensions the asymmetry operator will still be a pseudodifferential operator of order $-3$, as it is in the 3-dimensional case. However, a sufficient condition for a pseudodifferential operator to be of trace class is that its order be less than $-d$. This means that defining the regularised trace of the asymmetry operator becomes challenging: one would need to analyse singularities of the integral kernel and perform regularisation reducing the order of the singularity by $14k+$ 1. In other words, one would need to perform $14k+$ 1 steps of regularisation. Doing this in a geometrically invariant fashion will not be an easy task.

### v4

@@ 6.2 The order of the asymmetry operator
Definition 6.1 implies that, a priori, our asymmetry operator $A$ is a pseudodifferential operator of order 0. The goal of this subsection is to show that, as a result of cancellations in the symbol, the order of $A$ is in fact 3-3. This will be done in three steps.
@@
    First, we will show that $A$ is of order 2-2. This will follow easily from the facts already established earlier in the paper.
@@
    Next, we will show that, in normal coordinates, parallel transport does not contribute to homogeneous components of the symbol of $A$ of degree 2-2 and 3-3.
@@
    Finally, we will show that $A$ is of order 3-3. This is the subject of Theorem 6.6 which will come in the end of the subsection and is one of the main results of this paper.
@@
The asymmetry operator $A$ is a pseudodifferential operator of order 2-2.
@@
$p_{\pm}]_{\alpha}{}^{\beta}(x,\xi)\sim\sum_{j=0}^{+\infty}[(p_{\pm})_{-j}]_{\alpha}{}^{\beta}(x,\xi),\qquad[(p_{\pm})_{-k}]_{\alpha}{}^{\beta}(x,\lambda\xi)=\lambda^{-k}[(p_{\pm})_{-k}]_{\alpha}{}^{\beta}(x,\xi)\quad\forall\lambda>0,$ (6.5)
@@
The decomposition (6.6) is not invariant, it relies on our particular choice of local coordinates but it is convenient for our purposes. Namely, in what follows we will show that $A_{\mathrm{pt}}$ will not contribute to the principal symbol of order 3-3 of the asymmetry operator $A$ at the point $z$.
@@
In what follows, we will be using the expansion (F.8) for $Z_{\alpha}{}^{\beta}$.
The claim (6.9) for $j=0$ and $j=1$ follows immediately from (F.8).
@@
Upon integration by parts, we get from (F.8)
@@
where $|\cdot|$ is the Euclidean norm. By the symmetries of the Riemann tensor we have
$\operatorname{Riem}_{\alpha\mu\kappa\nu}=\frac{1}{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{Riem}}_{\alpha\mu\kappa\nu}+\operatorname{Riem}_{\kappa\nu\alpha\mu}).$ (6.11)
@@
The expansion (F.8) and formula (6.14) imply, upon integration by parts,
@@
The asymmetry operator $A$ is a pseudodifferential operator of order 3-3.
@@
The principal symbol $A_{\mathrm{prin},2}$ (that is, the principal symbol of the asymmetry operator $A$ viewed as an operator of order 2-2) should be proportional to
@@
$(a_{\mathrm{diag}})_{-2}](z,\xi)=[(p_{+}-p_{-})_{-2}]_{\alpha}{}^{\alpha}(z,\xi)=0.$ (6.22)
@@
Simplification 3. It is clear that $\operatorname{Ric}$ will appear $(a_{\mathrm{diag}})_{-2}](z,\xi)=[(p_{+}-p_{-})_{-2}]_{\alpha}{}^{\alpha}(0,\xi) in$ a linear fashion. This means that we can set
@@
  • one can scale $\xi$ so as to normalise it, using the homogeneity of $(a_{\mathrm{diag}})_{-2}](0,\xi)$.
@@
Simplification 5. In implementing the algorithm, expand all quantities at the $k$-th step in powers of $x$ and $\eta$ retaining only terms up to order 2k2-k, jointly in $x$ and $\eta$. Here we are assuming that $x$ and $\eta$ are of the same order.
Simplifications 1–5 reduce our algorithm to the manipulation of polynomials. Note that the vector spaces of polynomials in the variables $x^{\alpha}$ and $\eta_{\alpha}\,$, $\alpha=1,2,3$, of degree 0, 1 and 2 have dimension 1, 7 and 28 respectively [^4] The dimension of the vector space of polynomials of degree $k$ in $n$ variables is ${k+n}\choose{k}$}\,..
@@
$\operatorname{tr}(X_{+,2}-X_{-,2})_{\mathrm{prin}}](0,\xi_{0})=0.$ (6.29)
@@
where $(P_{\pm,1}^{2})_{-2}$ stands for the homogeneous component of degree 2-2 of the symbol of the operator $P_{\pm,1}^{2}\,$. To compute the latter, one needs to use the formula for the symbol of the composition of pseudodifferential operators: given two operators $A$ and $B$ with symbols $a$ and $b$, the symbol $\sigma_{BA}$ of their composition $BA$ is given by the formula
@@ 6.3 The principal symbol of the asymmetry operator
The principal symbol $A_{\mathrm{prin},3}$ (that is, the principal symbol of the asymmetry operator $A$ viewed as an operator of order 3-3) should be proportional to
@@
$(p_{+}-p_{-})_{-3}]_{\alpha}{}^{\alpha}(z,\xi)=-\frac{1}{2|\xi|^{5}}\varepsilon^{\alpha\beta\gamma}\,\nabla_{\alpha}\operatorname{Ric}_{\beta}{}^{\rho}(z)\,\xi_{\gamma}\xi_{\rho}\,.$ (6.32)
The strategy for proving (6.32) consists in implementing the algorithm from Proposition 5.3 in chosen normal coordinates up to order 3-3. Even with the help of modern computer algebra, doing this to the required accuracy is a highly nontrivial task. Therefore, as in the proof of Theorem 6.6, we describe below some helpful simplifications, which once again reduce all calculations to the manipulation of polynomials.
@@
Simplification 1. It is known, see [58, p. 211 formula (3.4)] or [55, formula (6)], that
@@
Simplification 5. In implementing the algorithm, expand all quantities at the $k$-th step in powers of $x$ and $\eta$ retaining only terms up order 3k3-k, jointly in $x$ and $\eta$. Here we are assuming that $x$ and $\eta$ are of the same order.
@@
$\operatorname{tr}(X_{+,3}-X_{-,3})_{\mathrm{prin}}](0,\xi_{0})=-\frac{1}{2}.$ (6.36)
@@
Since formula (1.13) is one of the key results of this paper, we provide in Appendix G an independent verification relying on formula (5.10).
@@ 7.1 Structure of the integral kernel near the diagonal
Now, the operator $A$ is not guaranteed to be of trace class, so one cannot apply formula (4.15) to it. The issue here is that in dimension 3 a sufficient condition for a pseudodifferential operator to be of trace class is that its order be strictly less than 3-3, whereas according to Theorem 1.4 for a generic Riemannian 3-manifold the order of $A$ is exactly 3-3. This means that we need to perform some sort of very basic regularisation in order to get over the finish line and define, in an invariant and geometrically meaningful manner, the trace of $A$ .
@@
This is an important property which will be used on two occasions in the subsequent arguments in this section. We will see that property (7.4) implies that the singularity of the integral kernel of our asymmetry operator $A$ is weaker than one would expect of a generic pseudodifferential operator of order 3-3. For a generic pseudodifferential operator of order 3-3 one would have a logarithm in the leading term of the singularity of the integral kernel, but there is no logarithm in (7.3).
@@
[…]
@@
$\int_{-\infty}^{+\infty}\frac{\cos\left(|y|\,\xi_{1}\right)}{(1+(\xi_{1})^{2})^{3/2}}\,d\xi_{1}=2|y|\,K_{1}(|y|)\,.$
The latter is known as Basset’s Integral [23, (10.32.11)]. ∎
@@
where $\gamma$ is the Euler–Mascheroni constant — see, e.g., [23, (10.31.11) and (10.25.2)]. By combining (7.9) and (7.6) one obtains (7.8). ∎
@@
In the previous section we have established that the asymmetry operator $A$ is a pseudodifferential operator of order 3-3 with principal symbol (1.13). Locally, the symbol $a(x,\xi)$ of $A$ and its integral kernel $\mathfrak{a}(x,y)$ are related via the (distributional) identity
@@
Since a pseudodifferential operator of order 4-4 on a 3-manifold has a continuous integral kernel, formulae (1.13), (1.10) and (7.11) imply that problem at hand reduces, in normal coordinates centred at $x$, to examining the quantity
$\frac{1}{16\pi^{3}}\varepsilon^{\alpha\beta\gamma}\,\nabla_{\alpha}\operatorname{Ric}_{\beta}{}^{\rho}(0)\,\int_{\mathbb{R}^{3}}e^{-iy^{\mu}\,\xi_{\mu}}\,\frac{\xi_{\gamma}\xi_{\rho}}{|\xi|^{5}}\,(1-\chi(|\xi|))\,\mathrm{d}\xi\,$,
@@
[…]
@@ 7.2 Local and global trace
[…]
@@
[…]
@@
$\frac{1}{4\pi r^{2}}\int_{\mathbb{S}_{r}(x)}\mathfrak{a}(x,y)\,dS_{y}=\frac{1}{4\pi r^{2}}\int_{\mathbb{S}_{r}(x)}\mathfrak{a}_{c}(x,y)\,\mathrm{d}S_{y}+\frac{c_{11}(x)+c_{22}(x)+c_{33}(x)}{3}\,.$ (7.16)
@@ 8 Challenges in higher dimensions
One could have also considered the case $d=4k+1$, $k=1,2,\ldots$, with $\operatorname{curl}$ acting in $\Omega^{2k}(M)$ in accordance with (1.5). However, in this case $\operatorname{curl}$ is formally anti-self-adjoint, the spectrum is purely imaginary and there is no spectral asymmetry [7, Theorem 3.2]. As our interest is in spectral asymmetry, we do not discuss the case $d=4k+1$. Note also that the spectral problem for $\operatorname{curl}$ in dimension $d=4k+1$ cannot be formulated in terms of the spectral theory of self-adjoint operators in real Hilbert spaces: one can introduce the factor i\,i\, in the right-hand side of (1.5) to make the operator formally self-adjoint (as was done in [7, Definition 2.2]), but the resulting operator will no longer be an operator acting in the vector space of real differential forms. The bottom line is that the two cases, $d=4k+3$ and $d=4k+1$, are fundamentally different. We stick with (8.1) because this gives the only genuine generalisation of the operator $\operatorname{curl}$ to higher dimensions.
@@
First challenge. We expect that in higher dimensions the asymmetry operator will still be a pseudodifferential operator of order 3-3, as it is in the 3-dimensional case. However, a sufficient condition for a pseudodifferential operator to be of trace class is that its order be less than d-d. This means that defining the regularised trace of the asymmetry operator becomes challenging: one would need to analyse singularities of the integral kernel and perform regularisation reducing the order of the singularity by $4k+1$. In other words, one would need to perform $4k+1$ steps of regularisation. Doing this in a geometrically invariant fashion will not be an easy task.